OpenStudy (anonymous):

Use the function rule to complete the table. –10x + y = 4 x –2 –1 0 1 2 y

OpenStudy (kaylaprincess):

-10(-2) + y = 4 -10(-1) + y = 4 -10(0) + y = 4 -10(1) + y = 4 -10(2) + y = 4 I started all the problems for you, if you need help solving them lemme know.

OpenStudy (kaylaprincess):

all you do is plug in x into the equation to get y :-)

OpenStudy (anonymous):

is it y = -10 x + y = 4

OpenStudy (kaylaprincess):

No, you must solve each equation to get y. I will do the first one for you, -10(-2) + y = 4 20 + y = 4 y = -20 + 4 y = - 16 So you would put -16 in the first part of the y chart.

OpenStudy (anonymous):

is the 3rd one 4

OpenStudy (anonymous):

@kaylaprincess is the 3rd one 4

OpenStudy (kaylaprincess):

correct!